OpenStudy (anonymous):

Solve the system of equations below. x-2y=3 2x-3=9 a. (3,9) b. (9,3) c. (-3,9) d. (-9,3)

OpenStudy (anonymous):

multiply the first one by (-2) and get: -2x+4y=-6

OpenStudy (anonymous):

right

OpenStudy (anonymous):

then, add the other one to it: -2x+4y=-6 2x-3y = 9 ---------- 0x+y = 3

OpenStudy (anonymous):

b is true

OpenStudy (anonymous):

@SwoleNationIzBack!! seriously what you say isnt gonna hurt me. so please leave me alone cause i have done nothing to you

OpenStudy (anonymous):

b is correct?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

just put each of them in the equations

OpenStudy (anonymous):

sure, we got this result from addition: -2x +4y =-6 2x -3y = 9 ------------- 0x+ y = 3 well, it states y=3. and you can put that in any equation to also get x! x-2y=3 x-2(3)=3 x-6=3 x = 9

OpenStudy (anonymous):

oh that clears things up. thank you!!!!